# Sensing atomic superfluid rotation beyond the standard quantum limit

### v1

@@ Abstract
Atomic superfluids formed using Bose-Einstein condensates (BECs) in a ring trap are currently being investigated in the context of superfluid hydrodynamics, quantum sensing and matter-wave interferometry. The characterization of the rotational properties of such superfluids is important, but can presently only be performed by using optical absorption imaging, which completely destroys the condensate. Recent studies have proposed coupling the ring BEC to optical cavity modes carrying orbital angular momentum to make minimally destructive measurements of the condensate rotation. The sensitivity of these proposals, however, is bounded below by the standard quantum limit set by the combination of laser shot noise and radiation pressure noise. In this work, we provide a theoretical framework that exploits the fact that the interaction between the scattered modes of the condensate and the light reduces to effective optomechanical equations of motion. We present a detailed theoretical analysis to demonstrate that the use of squeezed light and backaction evasion techniques allows the angular momentum of the condensate to be sensed with noise well below the standard quantum limit. Our proposal is relevant to atomtronics, quantum sensing and quantum information.
@@ I Introduction
Quantum sensing with highly controllable ultracold atoms [1, 2] is an actively developing field that provides new pathways to investigate foundational aspects of modern physics such as precise measurements of theoretically predicted fundamental constants [3, 4] and ultra-precise tracking of time [5, 6]. In recent years, there has been growing interest in the use of atomic superfluids such as Bose-Einstein condensates (BEC) confined inside a toroidal or ring trap for sensing [7, 8, 9]. Such configurations have been built using magnetic traps [10], time-averaged adiabatic potentials [11], painted dipole potentials [12], and intersecting “sheet” and “tube” laser beams [13]. Besides sensing, several fundamental and technologically relevant properties have been demonstrated ranging from persistent currents [10, 13] and quantum phase slips [14, 15] to atomic analogue of superconducting quantum interference device or SQUID [12] and atomtronic circuits [16, 17].
@@
More sophisticated techniques involving ring traps involve detecting the Doppler shift in phonon modes generated in the condensate [22] or by using an integrated atomic circuit, where information about the ring condensate is captured by studying the dynamics inside a tunnel-coupled rectilinear guide [23]. Recent studies [24, 25, 26] have proposed the use of quantum sensing techniques based on cavity optomechanics [27] to detect the rotational properties of superfluids in ring traps. The approach relies on the creation of sidemode excitations of the condensate due to interaction of the atoms with the orbital angular momentum (OAM) of the light [28] injected inside the cavity. While the above proposal allows for improvement in rotational sensitivity by a few orders of magnitude over other known techniques, it is still impacted by the noise imparted by the measurement setup, which restricts it to within the bounds of the standard quantum limit (SQL) [27, 29].
@@
The paper is arranged as follows. In Sec. II, the physical setup of a BEC confined in a toroidal trap, placed inside a cavity is presented. This is followed by Sec. III, where the theoretical framework for rotational sensing using squeezed OAM-carrying modes is presented. In Sec. IV, the backaction evasion method using a bichromatic driving field is introduced in the context of rotational sensing, with derivation of the equations of motion beyond the rotating wave approximation. Finally, the conclusions are presented in Sec. V.
@@ II Model
For sensing the BEC rotation, a laser beam of frequency $\omega$ is diffracted and recombined to form a superposed pair of beams with optical angular momenta ±$\pm l\hbar$ [51]. The cavity with resonant frequency $\omega_{a}$ and containing the trapped BEC, is then coherently driven by these OAM beams. Importantly, such a driving leads to the formation of an optical lattice around the axis of the cavity in the azimuthal direction (see Fig. 1). If the condensate rotating inside the cavity has a winding number $L_{p}$, a small number of the atoms will experience first-order Bragg scattering from the optical lattice [52] and attain the final winding numbers $L_{\pm}=L_{p}\pm 2l$, respectively. For a weak lattice [24], the number of atoms undergoing higher-order scattering will be negligible compared to number of atoms contained in the zeroth (unscattered) and first-order scattering sidemodes.
Figure 1: BEC in a toroidal ring trap inside an optical cavity. Superposed laser beams carrying OAM ±$\pm l$ (blue corkscrew wavefront and orange lines) are injected inside the cavity (blue lines). This creates an angular optical lattice (green regions) inside the ring trap (cyan toroidal shape) which contains the BEC. The input light is squeezed and shone from the bottom, while the output light after interaction is gathered from the top for homodyne detection and rotational sensing.
@@
where $\omega_{c(d)}=\left(L_{p}\pm 2l\right)^{2}\hbar/2I$ ($I$ is the moment of inertia of each atom about the center of the ring) and $a^{\dagger}$ and $a$ are the creation and annihilation operators, respectively, of the optical cavity field. The term ${\Delta}$ denotes the effective detuning between the driving and cavity frequency, $X_{c}=(c+c^{{\dagger}})/\sqrt{2}$ and $X_{d}=(d+d^{{\dagger}})/\sqrt{2}$ are the displacement operators of the atomic side modes, $G$ is the effective optomechanical interaction between the displacement $X_{c(d)}$ and the light intensity $a^{\dagger}a$, and $\eta$ is the driving amplitude of the input light. The final term $g\mathcal{Z}$ represents the effect of collisions between atoms in the BEC [24], where $g$ is the interaction between any two atoms.
@@ III Rotational sensing using squeezed light
The few-mode effective Hamiltonian in Eq. (2) describes the system consisting of the condensate with a winding number $L_{p}$ and interacting with the optical field carrying OAM ±$\pm l$$\hbar$. The dynamics of the system can be written in terms of the quantum Langevin equations,
$\ddot{X}_{c}+\gamma\dot{X}_{c}+\Omega_{c}^{2}X_{c}=-\tilde{\omega}_{c}Ga^{\dagger}a-\mathcal{A}X_{d}+\omega_{c}\epsilon_{c},$ (3)
$\ddot{X}_{d}+\gamma\dot{X}_{d}+\Omega_{d}^{2}X_{d}=-\tilde{\omega}_{d}Ga^{\dagger}a+\mathcal{A}X_{c}+\omega_{d}\epsilon_{d},$ (4)
@@
where $\Omega_{c(d)}^{2}=(\omega_{c(d)}+4gN)^{2}-4g^{2}N^{2}$ and $\mathcal{A}=2gN(\omega_{c}-\omega_{d})$. The damping and cavity loss rates are given by $\gamma$ and $\kappa$, respectively, and the shifted frequencies $\tilde{\omega}_{c(d)}=\omega_{c(d)}+2gN$. The driving term $\eta=\sqrt{\mathcal{P}_{\rm in}\kappa/\hbar\omega_{a}}$, where $\mathcal{P}_{\rm in}$ is the input laser power, and $\epsilon_{c(d),\rm in},a_{\rm in}$ are the input noise operators that take into account the thermal and quantum noise in the atomic sidemodes and the cavity modes, respectively.
In the Langevin formalism, the noise enters the equations in the form of correlation functions and the mean fluctuations are zero i.e., $\langle\epsilon_{c(d),\rm in}\rangle=\langle a_{\rm in}\rangle=0$. If the input cavity mode is squeezed, the general correlations of the noise operators are given by [27, 34],
@@
$\left\langle Q_{\text{in}}(\omega)P_{\text{in}}\left(\omega^{\prime}\right)\right\rangle=\left[1+\left(M_{r}^{*}-M_{r}\right)\right]\pi\delta\left(\omega+\omega^{\prime}\right),$
$\left\langle P_{\text{in}}(\omega)Q_{\text{in}}\left(\omega^{\prime}\right)\right\rangle=\left[-1+\left(M_{r}^{*}-M_{r}\right)\right]\pi\delta\left(\omega+\omega^{\prime}\right),$
$\left\langle\varepsilon_{k}(\omega)\varepsilon_{k}(\omega^{\prime})\right\rangle=\mathcal{B}_{k}\left(\coth\left[\frac{\hbar\omega_{k}}{2k_{B}T_{k}}\right]+1\right)\delta\left(\omega+\omega^{\prime}\right),$ (6)
where $Q_{\text{in}}=(a_{\text{in}}+a_{\text{in}}^{\dagger})/\sqrt{2}$ and $P_{\text{in}}=i(a_{\text{in}}^{\dagger}-a_{\text{in}})/\sqrt{2}$, are the quadrature operators of the incident light. Here, $\mathcal{B}=2\pi\gamma\omega/\omega_{k}$ and $T_{k}$ is the temperature of the BEC, where $k=\{c,d\}$. $N_{r}$ and $M_{r}$ are related to squeezing parameters $r$ and $\theta$ and are given by
@@
where, $N_{a}(\omega_{a})=1/({e^{\hbar\omega_{a}/k_{B}T}-1})$. $T$ is the ambient temperature. In the absence of squeezing, the terms $N_{r}$ and $M_{r}$ are identically zero, and the optical noise correlations reduce to $\langle a_{\rm in}(\omega)a_{\rm in}^{\dagger}(\omega^{\prime})\rangle=2\pi\delta\left(\omega+\omega^{\prime}\right)$. Moreover, $\hbar\omega_{a}\gg k_{B}T$, since thermal photons $N_{a}$ in the optical field are negligible.
For sensing the winding number $L_{p}$ of the BEC inside the ring trap, the central idea is to detect the phase quadrature of the cavity output field, which contains the two frequencies $\Omega_{c}$ and $\Omega_{d}$ [24]. A heuristic reading of Eqs. (3)-(4) in the absence of any noise and negligible radiation pressure shows that the sideband mode displacements $X_{c}$ and $X_{d}$ oscillate at frequencies $\Omega_{c}$ and $\Omega_{d}$, which in turn governs the optical field in Eq. (5), as the field operator $a$ is coupled to the two sideband modes. A homodyne detection of the output optical phase quadrature $P_{\rm out}(\omega)$ should therefore give a spectral density $S(\omega)$, which peaks at $\Omega_{c}$ and $\Omega_{d}$. Now, if $4g^{2}N^{2}\ll(\omega_{c(d)}+4gN)^{2}$, the spectral peaks are given by $\Omega_{c(d)}\approx\omega_{c(d)}+\Delta\omega_{\rm coll}$, where $\Delta\omega_{\rm coll}=2gN(2-gN)$. As such, by measuring the gap between the spectral peaks $\Omega_{c}-\Omega_{d}=\omega_{c}-\omega_{d}=4L_{p}l\hbar/I$, where $l$ and $I$ are already known, and the winding number $L_{p}$ of the BEC can be estimated. The objective now is to find the optimal conditions such that the measurement of output spectrum can be performed with minimal noise and high sensitivity. To achieve this, the equations of motion need to be solved to obtain the spectral density of the output optical field.
@@
$\delta u=\left[\delta Q,\delta P,\delta X_{c},\delta X_{d}\right]^{T}$ and $u_{\text{in}}=\left[Q_{\text{in}},P_{\text{in}},\epsilon_{c,\rm in},\epsilon_{d\rm in}\right]^{T}$ are vectors containing the fluctuation and input noise operators, respectively, such that $\delta Q=(\delta a+\delta a^{{\dagger}})/\sqrt{2}$,$\delta P=i(\delta a^{{\dagger}}-\delta a)/\sqrt{2}$, and $D={\rm diag}\left[\sqrt{\kappa},\sqrt{\kappa},\omega_{c},\omega_{d}\right]$ is a diagonal matrix with the system loss rates as elements. $\chi_{a}^{-1}(\omega)={\kappa}/{2}-i\omega$ and $\chi_{c(d)}^{-1}(\omega)=\Omega_{c(d)}^{2}-i\omega\gamma-\omega^{2}$, are susceptibility functions. The above matrix equation can be inverted to obtain the required analytic and numerical solutions for $\delta P(\omega)$ and $\delta Q(\omega)$. Now, using the input-output formalism [53, 44], the output optical quadrature can be found from the relation, $\delta K_{\rm out}(\omega)=\sqrt{\kappa}\delta K(\omega)-\delta K_{\rm in}(\omega)$, for $K=\{Q,P\}$. Note that the prefix $\delta$ in front of input and output terms is dropped henceforth, for compactness. The analytical expressions for the coefficients for these terms are provided in Appendix A.
Figure 2: Variation of noise spectral density, sensitivity with frequency. (a) The plot shows the output optical noise spectral density for unsqueezed ($r=0$, solid blue line) and squeezed ($r=2,\theta=\pi$, dashed orange line) input, both without inter-atomic collisions ($g=0$). For cases where inter-atomic interactions are included ($g\neq 0$) is indicated with green dash-dot line for $r=0$ and with red dotted curve for $r=2,\theta=\pi$. For squeezing angle $\theta=\pi$, the noise spectrum falls below the shot noise level (black-dotted horizontal line). b) Shows the variation of sensitivity for the same set of parameters above. The input power $\mathcal{P}_{\rm in}=12.4$ fW and the detection angle $\phi=\pi/2$. The parameters $N=10^{4}$, $R=12~{}\mu$m, $gN\approx 14$ Hz, $I/\hbar=0.0505$ Hz, and $G/2\pi=7.5$ KHz, correspond primarily to a condensate of sodium atoms [24]. Other parameters are $\omega_{a}/2\pi=10^{2}$ THz, $\omega_{r}/2\pi=\omega_{z}/2\pi=42$ Hz, $\kappa/2\pi=2$ MHz, $\gamma/2\pi=0.8$ Hz, $\Delta^{\prime}=0$, and $T_{c,d}=20$ nK.
@@
where $X_{\rm in},Y_{\rm in}\in\{Q_{\rm in},P_{\rm in},\epsilon_{c},\epsilon_{d}\}$. The calculations for the detection angle $\phi$ that optimize the noise spectrum and sensitivity under different conditions are provided in Appendix B.
@@
$\zeta(\omega)=\dfrac{S(\omega)}{\partial S(\omega)/\partial\Lambda}\sqrt{t_{\rm{meas}}},$
$\zeta_{\rm{opt}}=\zeta(\omega=\omega_{\rm{opt}});~{}\omega_{\rm{opt}}:~{}\left[\dfrac{\partial\zeta}{\partial\omega}\right]_{\omega=\omega_{\rm{opt}}}=~{}0,$ (13)
where $t_{\rm meas}\approx\kappa/8a_{s}^{2}G^{2}$ is the measurement time in the bad cavity regime [27] and $S(\omega)$ responds linearly with condensate angular momentum, $\Lambda=L_{p}\hbar$. The optimal sensitivity $\zeta_{\rm{opt}}$ corresponds to the minima of the function $\zeta(\omega)$ at frequency $\omega_{\rm opt}$.
Figure 2 shows the variation of $S^{\rm out}_{P,P}(\omega)$ and $\zeta(\omega)$ with frequency $\omega$, and studies the effect of introducing squeezed input and atomic collisions in the condensate by solving Eqs. (3)-(5). The winding number of the input light and the BEC are taken as $l=10$ and $L_{p}=1$, respectively, which results in peaks of the output spectrum at $\Omega_{c(d)}/2\pi$ at 569 and 695 Hz, respectively, without collisions and 624 and 750 Hz, respectively, with collisions, when $gN=14$ Hz. The optimal sensitivity $\zeta_{\rm opt}$ also occurs at $\omega_{\rm opt}\approx\Omega_{c(d)}$. The plots in Fig. 2 show that the presence of inter-atomic interactions $g$, simply shifts the frequencies of the spectral peaks and also the optimal sensitivity, for both squeezed and unsqueezed light (cf. [24]). However, for squeezed light with amplitude $r=2$ and angle $\theta=\pi$, the spectral density drops below the shot noise level, which is an indicator that the noise in the spectrum can be lowered below the quantum limi. For detection angle $\phi=\pi/2$ and squeezing angle $\theta=\pi$, the optimal sensitivity $\zeta_{\rm opt}$ of measurement is not significantly enhanced over the unsqueezed case.

### v2

@@ Abstract
Atomic superfluids formed using Bose-Einstein condensates (BECs) in a ring trap are currently investigated in the context of superfluid hydrodynamics, quantum sensing and matter-wave interferometry. The characterization of the rotational properties of such superfluids is important, but can presently only be performed by using optical absorption imaging, which completely destroys the condensate. Recent studies have proposed coupling the ring BEC to optical cavity modes carrying orbital angular momentum to make minimally destructive measurements of the condensate rotation. The sensitivity of these proposals, however, is bounded below by the standard quantum limit set by the combination of laser shot noise and radiation pressure noise. In this work, we provide a theoretical framework that exploits the fact that the interaction between the scattered modes of the condensate and the light reduces to effective optomechanical equations of motion. We present a detailed theoretical analysis to demonstrate that the use of squeezed light and backaction evasion techniques allows the angular momentum of the condensate to be sensed with noise well below the standard quantum limit. Our proposal is relevant to atomtronics, quantum sensing and quantum information.
@@ I Introduction
Quantum sensing with highly controllable ultracold atoms [1, 2] is an actively developing field that provides new pathways to investigate foundational aspects of modern physics such as precise measurements of theoretically predicted fundamental constants [3, 4] and ultra-precise tracking of time [5, 6]. In recent years, there has been growing interest in the use of atomic superfluids such as Bose-Einstein condensates (BEC) confined inside a toroidal or ring trap for sensing [7, 8, 9]. Such configurations have been built using magnetic traps [10], time-averaged adiabatic potentials [11], painted dipole potentials [12], and intersecting “sheet” and “tube” laser beams [13]. Besides sensing, several fundamental [14, 15, 16] and technologically relevant properties have been demonstrated ranging from persistent currents [10, 13] and quantum phase slips [17, 18] to atomic analogs of superconducting quantum interference devices [12] and atomtronic circuits [19, 20].
@@
More sophisticated techniques involving ring traps involve detecting the Doppler shift in phonon modes generated in the condensate [24] or using an integrated atomic circuit, where information about the ring condensate is captured by studying the dynamics inside a tunnel-coupled rectilinear guide [25]. Recent studies [26, 27, 28] have proposed the use of quantum sensing techniques based on cavity optomechanics [29] to detect the rotational properties of superfluids in ring traps. The approach relies on the creation of sidemode excitations of the condensate due to interaction of the atoms with the orbital angular momentum (OAM) of the light [30] injected inside the cavity. While the above proposal allows for improvement in rotational sensitivity by a few orders of magnitude over other known techniques, it is still impacted by the noise imparted by the measurement setup, which restricts it to within the bounds of the standard quantum limit (SQL) [29, 31].
@@
The paper is arranged as follows. In Sec. II, the physical setup of a BEC confined in a toroidal trap, placed inside a cavity is presented. This is followed by Sec. III, where the theoretical framework for rotational sensing using squeezed OAM-carrying modes is presented. In Sec. IV, the backaction evasion method using a bichromatic driving field is introduced in the context of rotational sensing, with derivation of the equations of motion beyond the rotating wave approximation. A summary and conclusions are presented in Sec. V.
@@ II Model
For sensing the BEC rotation, a laser beam of frequency $\omega$ is diffracted and recombined to form a superposed pair of beams with optical angular momenta ±$\pm l\hbar$ [53]. The cavity with resonant frequency $\omega_{a}$ and containing the trapped BEC, is then coherently driven by these OAM beams. Importantly, such a driving leads to the formation of an optical lattice around the axis of the cavity in the azimuthal direction (see Fig. 1). If the condensate rotating inside the cavity has a winding number $L_{p}$, a small number of the atoms will experience first-order Bragg scattering from the optical lattice [54] and attain the final winding numbers $L_{\pm}=L_{p}\pm 2l$, respectively. For a weak lattice [26], the number of atoms undergoing higher-order scattering will be negligible compared to number of atoms contained in the zeroth (unscattered) and first-order scattering sidemodes. We note that even the first order scattered sidemodes contain a very small fraction of the total atoms in the condensate. As such, the sidemodes have a negligible chemical potential and the temperature is not significantly affected due to scattering. This is in contrast to cases where splitting of BEC occurs to form new condensed modes [55].
Figure 1: BEC in a toroidal ring trap inside an optical cavity. Superposed laser beams carrying OAM ±$\pm l\hbar$ (blue corkscrew wavefront and orange lines) are injected inside the cavity (blue lines). This creates an angular optical lattice (green regions) inside the ring trap (cyan toroidal shape) which contains the BEC. The input light is squeezed and shone from the bottom, while the output light after interaction is gathered from the top for homodyne detection and rotational sensing.
@@
where $\omega_{c~{}(d)}=\left(L_{p}\pm 2l\right)^{2}\hbar/2I$ ($I$ is the moment of inertia of each atom about the center of the ring) and $a^{\dagger}$ and $a$ are the creation and annihilation operators, respectively, of the optical cavity field. The term ${\Delta}$ denotes the effective detuning between the driving and cavity frequency. The field injected in the cavity is a single mode superposition of frequency degenerate OAM light, |$\ket{+l\hbar}+e^{i\varphi_{0}}\ket{-l\hbar}$, where the relative phase $\varphi_{0}$ can be fixed by the experimentalist. The light intensity in the corresponding optical lattice varies as $\cos^{2}(l\varphi-\varphi_{0}/2)$, where $\varphi$ is the azimuthal degree of freedom. The scattering with atoms in the BEC only transforms |±$\ket{\pm l\hbar}\rightarrow\ket{\mp l\hbar}$, and the optical lattice remains unchanged. This implies no new optical modes are populated. In this work we consider $\varphi_{0}=0$, which gives us a symmetric superposition of |±$\ket{\pm l\hbar}$ (see Appendix A for further details). The displacement operators of the atomic side modes are given by $X_{c}=(c+c^{{\dagger}})/\sqrt{2}$ and $X_{d}=(d+d^{{\dagger}})/\sqrt{2}$, where the effective optomechanical interaction between the displacement $X_{c(d)}$ and the optical lattice is given by $G$. $\eta$ is the driving amplitude of the input beam. The final term $g\mathcal{Z}$ represents the effect of collisions between atoms in the BEC, where $g$ is the interaction between any two atoms. The term $G$ and operator $\mathcal{Z}$ are derived in Ref. [26], with $G=g_{a}^{2}\sqrt{N}/2\sqrt{2}\Delta_{a}$, where $g_{a}$ is the coupling strength of a photon with a single atom, $\Delta_{a}$ is the detuning of input beam with the atomic transition level. $\mathcal{Z}$ contains collision operators up to order $N$, which significantly contributes to the dynamics.
@@ III Rotational sensing using squeezed light
The few-mode effective Hamiltonian in Eq. (2) describes the system consisting of the condensate with a winding number $L_{p}$ and interacting with the optical field carrying OAM ±$\pm l\hbar$. The dynamics of the system can be written in the form of the quantum Langevin equations [29, 56]. They can be obtained by writing the Heisenberg equations of motion for the cavity and the sidemode operators. Below, we express the equations in terms of only the cavity and position quadrature of sidemodes, which resemble a pair of damped oscillators, driven by photons and stochastic noise operators,
$\ddot{X}_{c}+\gamma\dot{X}_{c}+\Omega_{c}^{2}X_{c}=-\tilde{\omega}_{c}Ga^{\dagger}a-\mathcal{A}X_{d}+\omega_{c}\epsilon_{c,\mathrm{in},}$ (3)
$\ddot{X}_{d}+\gamma\dot{X}_{d}+\Omega_{d}^{2}X_{d}=-\tilde{\omega}_{d}Ga^{\dagger}a+\mathcal{A}X_{c}+\omega_{d}\epsilon_{d,\mathrm{in}},$ (4)
@@
where $\omega_{c~{}(d)}^{2}=(\omega_{c~{}(d)}+4gN)^{2}-4g^{2}N^{2}$ and $\mathcal{A}=2gN(\omega_{c}-\omega_{d})$ are the modified sidemode frequencies and coupling between them. The contributions from the interatomic collisions with strength $g$ arise from the operator $\mathcal{Z}$ in Eq. (2)[26]. The damping and cavity loss rates are given by $\gamma$ and $\kappa$, respectively, and the shifted frequencies $\tilde{\omega}_{c(d)}=\omega_{c~{}(d)}+2gN$. The driving term $\eta=\sqrt{\mathcal{P}_{\rm in}\kappa/\hbar\omega_{a}}$, where $\mathcal{P}_{\rm in}$ is the input laser power, and $\epsilon_{c(d),\rm in},a_{\rm in}$ are the input noise operators that take into account the thermal and quantum noise in the atomic sidemodes and the cavity modes, respectively.
In the Langevin formalism [29, 56], the noise enters the equations in the form of correlation functions, where the mean fluctuations are zero i.e., $\langle\epsilon_{c(d),\rm in}\rangle=\langle a_{\rm in}\rangle=0$. We introduce squeezing in the optical modes using the input-output formalism, which allows for modeling quantum fluctuations injected into the cavity, in addition to any laser driven coherent state. As such, squeezed vacuum fluctuations $a_{\rm in}$ are added on top of the coherent cavity state $a$, which can be experimentally realised using a combination of beamsplitters and an optical parametric oscillator [57, 58]. Now, if the input cavity mode is squeezed, the general correlations of the noise operators are given by [29, 36],
@@
$\left\langle Q_{\text{in}}(\omega)P_{\text{in}}\left(\omega^{\prime}\right)\right\rangle=i\left[1+\left(M_{r}^{*}-M_{r}\right)\right]\pi\delta\left(\omega+\omega^{\prime}\right),$
$\left\langle P_{\text{in}}(\omega)Q_{\text{in}}\left(\omega^{\prime}\right)\right\rangle=i\left[-1+\left(M_{r}^{*}-M_{r}\right)\right]\pi\delta\left(\omega+\omega^{\prime}\right),$
$\left\langle\epsilon_{k,\rm{in}}(\omega)\epsilon_{k,\rm{in}}(\omega^{\prime})\right\rangle=\mathcal{B}_{k}\left(\coth\left[\frac{\hbar\omega_{k}}{2k_{B}T_{k}}\right]+1\right)\delta\left(\omega+\omega^{\prime}\right),$ (6)
where $Q_{\text{in}}=(a_{\text{in}}+a_{\text{in}}^{\dagger})/\sqrt{2}$ and $P_{\text{in}}=i(a_{\text{in}}^{\dagger}-a_{\text{in}})/\sqrt{2}$, are the quadrature operators of the incident light. Here, $\mathcal{B}_{k}=2\pi\gamma\omega/\omega_{k}$ and $T_{k}$ is the temperature of the BEC, where $k=\{c,d\}$. $N_{r}$ and $M_{r}$ are related to squeezing parameters $r$ and $\theta$ and are given by
@@
where $N_{a}(\omega_{a})$ is the number of thermal photons, which we have neglected as cavity frequency is in the optical regime. Moreover, in the absence of squeezing, the terms $N_{r}$ and $M_{r}$ are identically zero, and the optical noise correlations reduce to $\langle a_{\rm in}(\omega)a_{\rm in}^{\dagger}(\omega^{\prime})\rangle=2\pi\delta\left(\omega+\omega^{\prime}\right)$.
For sensing the winding number $L_{p}$ of the BEC inside the ring trap, the central idea is to detect the phase quadrature of the cavity output field, which contains the two frequencies $\Omega_{c}$ and $\Omega_{d}$ [26]. A heuristic reading of Eqs. (3)-(4) in the absence of any noise and negligible radiation pressure shows that the sideband mode displacements $X_{c}$ and $X_{d}$ oscillate at frequencies $\Omega_{c}$ and $\Omega_{d}$, which in turn governs the optical field in Eq. (5), as the field operator $a$ is coupled to the two sideband modes. A homodyne detection of the output optical phase quadrature $P_{\rm out}(\omega)$ should therefore give a spectral density $S(\omega)$, which peaks at $\Omega_{c}$ and $\Omega_{d}$. Now, if $4g^{2}N^{2}\ll(\omega_{c~{}(d)}+4gN)^{2}$, the spectral peaks are given by $\omega_{c~{}(d)}\approx\omega_{c~{}(d)}+\Delta\omega_{\rm coll}$, where $\Delta\omega_{\rm coll}=2gN(2-gN/\omega_{c~{}(d)})$. As such, by measuring the gap between the spectral peaks $\Omega_{c}-\Omega_{d}=\omega_{c}-\omega_{d}=4L_{p}l\hbar/I$, where $l$ and $I$ are already known, and the winding number $L_{p}$ of the BEC can be estimated. The objective now is to find the optimal conditions such that measurement of output spectrum can be performed with minimal noise and high sensitivity. To achieve this, the equations of motion need to be solved to obtain the spectral density of the output optical field.
@@
$\delta u=\left[\delta Q,\delta P,\delta X_{c},\delta X_{d}\right]^{T}$ and $u_{\text{in}}=\left[Q_{\text{in}},P_{\text{in}},\epsilon_{c,\rm in},\epsilon_{d,\rm in}\right]^{T}$ are vectors containing the fluctuation and input noise operators, respectively, such that $\delta Q=(\delta a+\delta a^{{\dagger}})/\sqrt{2}$,$\delta P=i(\delta a^{{\dagger}}-\delta a)/\sqrt{2}$, and $D={\rm diag}\left[\sqrt{\kappa},\sqrt{\kappa},\omega_{c},\omega_{d}\right]$ is a diagonal matrix with the system loss rates as elements. $\chi_{a}^{-1}(\omega)={\kappa}/{2}-i\omega$ and $\chi_{c(d)}^{-1}(\omega)=\omega_{c~{}(d)}^{2}-i\omega\gamma-\omega^{2}$, are susceptibility functions. The above matrix equation can be inverted to obtain the required analytic and numerical solutions for $\delta P(\omega)$ and $\delta Q(\omega)$. Now, using the input-output formalism [59, 46], the output optical quadrature can be found from the relation, $\delta K_{\rm out}(\omega)=\sqrt{\kappa}\delta K(\omega)-\delta K_{\rm in}(\omega)$, for $K=\{Q,P\}$. Note that the prefix $\delta$ in front of input and output terms is dropped henceforth, for compactness. The analytical expressions for these coefficients are provided in Appendix B.
Figure 2: Variation of spectral density and sensitivity with frequency. (a) The plot shows the output optical noise spectral density for unsqueezed ($r=0$, solid blue line) and squeezed ($r=2,\theta=\pi$, dashed orange line) input, without inter-atomic collisions ($g=0$). Cases with inter-atomic interactions ($g\neq 0$) are indicated with green dash-dot line for $r=0$ and with red dotted curve for $r=2,\theta=\pi$. For squeezing angle $\theta=\pi$, the noise spectrum falls below the shot noise level (black-dotted horizontal line). b) Shows the variation of sensitivity for the same set of parameters above. The input power $\mathcal{P}_{\rm in}=12.4$ fW and the detection angle $\phi=\pi/2$. The parameters $N=10^{4}$, $R=12~{}\mu$m, $gN\approx 14$ Hz, $I/\hbar=0.0505$ Hz, and $G/2\pi=7.5$ kHz, correspond primarily to a condensate of sodium atoms [26]. Other parameters are $\omega_{a}/2\pi=10^{3}$ THz, $\omega_{r}/2\pi=\omega_{z}/2\pi=42$ Hz, $\kappa/2\pi=2$ MHz, $\gamma/2\pi=0.8$ Hz, $\Delta^{\prime}=0$, $T_{c,d}=20$ nK, $l=10$, and $L_{p}=1$.
@@
where $X_{\rm in},Y_{\rm in}\in\{Q_{\rm in},P_{\rm in},\epsilon_{c},\epsilon_{d}\}$. The calculations for the detection angle $\phi$ that optimize the noise spectrum and sensitivity under different conditions are provided in Appendix C.
@@
$\zeta(\omega)=\dfrac{S(\omega)}{\partial S(\omega)/\partial\Lambda}\sqrt{t_{\rm{meas}}},~{}\zeta_{\rm{opt}}=\zeta(\omega=\omega_{\rm{opt}}),$ (13)
where $t_{\rm meas}\approx\kappa/8a_{s}^{2}G^{2}$ is the measurement time in the bad cavity regime [29] and $\omega_{\rm opt}$ is defined as $\left(\partial\zeta/\partial\omega\right)_{\omega=\omega_{\rm{opt}}}=0$. The optimal sensitivity $\zeta_{\rm{opt}}$ corresponds to the minima of the function $\zeta(\omega)$ at frequency $\omega_{\rm opt}$. The motive here is to find an optimal frequency where the change in spectrum $S(\omega)$ is maximum, which is often relevant in experimental implementations using a lock-in detection scheme [60]. Moreover, $S(\omega)$ responds linearly with angular momentum $\Lambda=L_{p}\hbar$ at values of $L_{p}$ relevant to our analysis, such that $S(\omega)\left[\frac{\Delta S(\omega)}{\Delta\Lambda}\right]^{-1}\simeq S(\omega)\left[\frac{\partial S(\omega)}{\partial\Lambda}\right]^{-1}$, which is useful for numerical estimation of sensitivity.
Figure 2 shows the variation of $S^{\rm out}_{P,P}(\omega)$ and $\zeta(\omega)$ with frequency $\omega$, and studies the effect of introducing squeezed input and atomic collisions in the condensate by solving Eqs. (3)-(5). The winding number of the input light and the BEC are taken as $l=10$ and $L_{p}=1$, respectively, which results in peaks of the output spectrum at $\omega_{c~{}(d)}/2\pi$ at 569 and 695 Hz, respectively, without collisions and 624 and 750 Hz, respectively, with collisions, when $gN=14$ Hz. The optimal sensitivity $\zeta_{\rm opt}$ also occurs at $\omega_{\rm opt}\approx\omega_{c~{}(d)}$. The plots in Fig. 2 show that the presence of inter-atomic interactions $g$, simply shifts the frequencies of the spectral peaks and also the optimal sensitivity, for both squeezed and unsqueezed light (cf. [26]). However, for squeezed light with amplitude $r=2$ and angle $\theta=\pi$, the spectral density drops below the shot noise level, which is an indicator that the noise in the spectrum can be lowered below the quantum limit. For detection angle $\phi=\pi/2$ and squeezing angle $\theta=\pi$, the optimal sensitivity $\zeta_{\rm opt}$ of measurement is not significantly enhanced over the unsqueezed case.